OpenStudy (anonymous):

Simplify (4x^2 + 3x – 2) – (2x^2 - 4x + 4)

OpenStudy (anonymous):

2x^2+7x-6

OpenStudy (anonymous):

would you like an explanation or you're good on that?

OpenStudy (anonymous):

@trueseminole Yea

OpenStudy (anonymous):

okay 4x^2 - 2x^2 = 2x^2 3x+ 4x = 7x 2 + 4 = 6

OpenStudy (anonymous):

@trueseminole thanks so much! the test i am taking is the difference in me getting a B or C in the class so it is important for me to do well on it!

OpenStudy (anonymous):

well good luck